# How Linguistics Learned to Stop Worrying and Love the Language Models

### v1

@@ 3.3.2 The statistical tradition of linguistics
Perhaps the most well-known such contribution is distributional semantics: the idea that the semantics of a word is related to the distribution over contexts in which that word appears, often cited to Firth, (1957, p. 11), and developed more systematically by Harris, (1954). This idea originates out of the school of structuralist linguistics (Saussure,, 1916; Bloomfield,, 1926). Structuralist linguistics had as one theoretical aim the development of discovery procedures, which were formal procedures which could be applied to bodies of text in order to discover (and define) linguistic structures (Harris,, 1951). The most well-developed of these discovery procedures were statistical in nature. For example, Harris, (1955) developed a theory of words and morphemes based on statistical co-occurrence patterns, including a procedure for discovering morpheme boundaries by effectively calculating transitional probabilities, an idea taken up again much later in the psycholinguistics of language learning (Saffran et al.,, 1996), and closely related to tokenization methods such as byte-pair encoding (Shibata et al.,, 1999; Tanaka-Ishii,, 2021, Ch. 11). At issue was the relationship between grammatical structure and the observable statistical structure of a corpus of language—one of the core questions that linguists working on language models are interested in again today.
@@
Nevertheless, the statistical analysis of language continued under the heading of usage-based approaches (Bybee and Hopper,, 2001), in which the key to language is not a set of underlying formal rules but emergent properties based on the statistics and dynamics of language use (from which rules or rule-like behavior might emerge). Traditions in typological syntax (Greenberg, 1963a, ; Dryer,, 1992), functionalist syntax (Comrie,, 1989; Keenan and Comrie,, 1977), construction grammar (Croft,, 2001; Goldberg,, 2009), probabilistic modeling (Bresnan et al.,, 2001, 2007; Christiansen and Chater,, 2016), and evolutionary linguistics (Kirby and Hurford,, 2002) have all carried the banner of a thoroughgoingly statistical approach to language.
@@
In fact, these statistical traditions of linguistics played a catalyzing role in the birth of neural networks and LMs. While it is sometimes claimed that LMs are purely a triumph of engineering, the connectionist framework from which they emerged overlapped with the statistical tradition in linguistics, often in conflict with the anti-statistical tradition. A major locus of this early work using neural networks was about the English past tense, with Rumelhart and McClelland, (1987) developing a neural model for forming the past tense from present tense words, triggering much debate (Pinker and Prince,, 1988). Elman, 1990b, a Linguistics PhD, invented the recurrent neural network (RNN) architecture () in order to solve the problem of finding linguistic structure in time, a problem motivated by findings in linguistics and psycholinguistics (Frazier and Fodor,, 1978). This basic architecture still underlies many language models (Feng et al.,, 2024). Chris Manning, a pioneer in neural and probabilistic models, was trained as a linguist (Manning,, 1995), and argued for a statistical approach to syntax (Manning,, 2003), before going on to develop influential neural methods for performing linguistic tasks (e.g., Pennington et al.,, 2014).
@@
The history of the field and the dominance of the generative tradition made it such that these figures are sometimes seen as ‘not linguists’. But the traditions they emerged from were fundamentally concerned with questions about human language and cognition, so we believe they do belong under the heading of linguistics. To be clear, we do not think that the success of modern language models unequivocally supports or refutes any particular intellectual tradition in linguistics. But we do think that the narrow and exclusive theoretical focus of generative linguistics, coupled with its relative dominance within American linguistics departments throughout the late 20th century, caused tragic missed connections in intellectual history.
@@
Progress in language science will not come from hyperfocus on any one goal, nor from any one theory of what language is, nor from one framework for understanding linguistic phenomena. Rather, the science of language needs to approach as many questions as possible from as many angles as possible. In the current moment, that means leveraging the explosion of intellectual creativity springing forth around language models. Just as the concept of entropy arose from ad-hoc analysis of complex machines and ended up revolutionizing fundamental physics, it is possible and even likely that ideas based on language models will revolutionize linguistics.
@@ 4 Where does that leave the science of language?
Where do we go from here? What is the status of the formal structures discovered through linguistic analysis, if they don’t have to be innately latent in the human brain? How can we explain why language is the way it is? What approach toward language is most productive in an era when we have huge neural networks that seem to be able to produce and comprehend it?
We believe that linguistic structure is as real as it ever was, and that the LM revolution presents us with exciting new ideas and methods for studying language. Taking a functional and information-theoretic approach brings these possibilities into focus. We also consider ways in which LMs might give us reason to update some of our linguistic theories to be more gradient and usage-based, by serving as a proof-of-concept system.
From an information-theoretic perspective, as we will argue in Section 4.1, linguistic structure is real because it is an abstraction that allows us to compress the myriad data of linguistic experience in a way that lets us predict, understand, and organize that experience. Even if it is not hardwired in the brain, or anywhere, linguistic structure is a necessary concept in the sense that without it we would have no realistic hope of coherently describing linguistic data in a way that humans can understand—and that goes for language models as well as human linguistic behavior. At the same time, we think that language models can help us study what kinds of theories successfully do explain and help us understand language data and linguistic behavior.
As for answering the question of why language is the way it is, as we argue in Section 4.4, the information-theoretic perspective invites us to consider language at the functional level, where we find that human language processing and language models are more similar than one might have expected. Thus, language models can inspire explanatory theories of language in the same way that convolutional networks inspired explanatory theories of vision (see Section 3.1), and indeed we see that the neural network literature spawns multiple exciting explanatory hypotheses for human language. Considering the broader function of language—allowing humans to communicate with each other subject to constraints on how comprehension and production work—brings language models to the forefront of linguistics as indispensable tools.
@@ 4.1 Linguistic structure is real
If a general-purpose sequence model does learn language, but it does not build in the hierarchical representations common to models of linguistic structure, then what is the status of those representations and models? We believe that linguistic structure is nevertheless real in the sense of Dennett, (1989). Linguistic structure provides a compressed and useful representation of language. To explain or describe linguistic phenomena without reference to linguistic structure would be hopelessly complex, just like Marr,’s example of trying to explain bird wings by studying only feathers.
Consider a phenomenon like subject–verb number agreement. Modern language models seem to be reasonably proficient at it, even when they are trained on corpora with no examples where a verb must agree with a non-local noun (Patil et al.,, 2024). As we argued earlier, we think this happens because they are meaningfully learning grammatical structure. But consider a thought experiment: imagine that language models learned to do subject–verb agreement through some complicated and esoteric statistical process, picking up on subtle cues in data that we have never identified or thought of. Even though this complex process doesn’t seem to straightforwardly map onto our linguistic account, we find that the model is behaviorally perfect at subject–verb agreement.
@@
Hypothetically, given that a system learned this mechanism, it might be worth investigating whether people are actually implementing some version of it, treating language models as hypothesis generators (e.g., Lakretz et al.,, 2021; Portelance and Jasbi,, 2023). If it turns out that there are in fact messy and heterogenous processes that give rise to rule-like behavior even in humans, one reaction might be to conclude that linguistic theory was wrong or misguided—that it seemed like there was such a thing as “grammatical subject”, but actually the underlying reality was more complicated and now we can do away with grammatical subject and agreement rules and replace them with the messy, complicated neural system.
@@
This hyper-reductive conclusion would not be productive. A theory of language which points to complex neural circuits for explaining the data of subject–verb agreement, without understanding something about the concept of grammatical subject, would be like an example from Dennett, (1989, Ch. 2): imagine aliens who can perfectly predict human behavior using atomic physics (that clump of human atoms will move towards that coffee shop) but without understanding human beliefs and desires (that the human wants coffee and believes that the coffee shop will provide it). Lacking concepts of humans’ attitudes, the aliens’ theory of human behavior is perhaps perfectly accurate and reductive, but comes “at enormous cost of cumbersomeness, lack of generality, and unwanted detail” (Dennett,, 1990, p. 189). Dennett, concludes that the concept of belief is a real pattern in that it provides a (potentially not 100% accurate) abstraction that supports prediction and counterfactual reasoning based on coarse-grained data. The idea is that any abstraction that enables simple prediction, compression, and causal modeling in this way is a real pattern, fully deserving to be called real even if there is a more accurate, more reductive theory that lacks the abstraction. Similarly, even if it turned out we could explain linguistic behavior without recourse to anything recognizable as linguistic structure, we believe it would come at the same cost of cumbersomeness, lack of generality, and unwanted detail.
@@
Figure 3: Schematic for linguistic structure as a real pattern, a (potentially leaky) abstraction that supports prediction and counterfactual reasoning based on coarse-grained data. In neural networks (in silico), the complex patterns of weights and activations may be abstracted into interpretable circuits that approximately compute linguistically meaningful features like whether a word is singular and a grammatical subject. In turn, those interpretable circuits may be interpretable in terms of a larger abstraction of a coherent linguistic structure. In humans (in vivo), the complex biological and social processes associated with language learning, use, and knowledge, may be abstracted into psycholinguistic theories describing operations on mental representations. In turn, the behaviors of these mental representations may be interpretable in terms of coherent linguistic structure. Thus linguistic structure is a real pattern at the highest level of abstraction, which allows us to understand language as it is implemented by a human or by a neural network.
Thus we join Nefdt, (2023) in seeing linguistic structure and theory as giving us real patterns—a position consistent with the position laid out in Smolensky, (1988) that “Parallel connectionist memory and linguistic processes are hypothesized to give rise to processes that are describable at a higher level as sequential rule application.” And we join Dennett, in his conviction that real patterns are, well, real. That is, we are not eliminativists about grammatical structure, nor do we believe that this kind of eliminativism would be scientifically productive. Our view is summarized in Figure 3.
Versions of our view have been a common refrain from all sides. Rumelhart and McClelland, 1986a write: “we suggest that the mechanisms that process language and make judgements of grammaticality are constructed in such a way that their performance is characterizable by rules, but that the rules themselves are not written in explicit form anywhere in the mechanism”. Some in the generative tradition seem to express the same view, like Adger, (2022): “Generative grammarians think of the posits of their theories in much the same way as physicists think of quarks or particles in wave form: They are the best explanations of the phenomena, though we do not know exactly what they correspond to in the world we can currently observe.” While there are many ways for a particular theory to be implemented computationally (Stabler,, 1983), LMs are a promising avenue for figuring out what these explanations do, in fact, correspond to at a lower level— insofar as they are successful (see earlier discussion of the Contravariance Principle) and insofar as linguistic structure provides efficient, useful representations of linguistic data.
@@
Indeed, the real patterns framework can help explain the success of language models. It is exactly because linguistic structure is a real pattern that models are able to learn linguistic structure from data. Traditional language models are trained to predict tokens in context. From an information-theoretic perspective, this means they are finding simple representations for linguistic input (Tishby et al.,, 1999; Chater and Vitányi,, 2007; Tishby and Zaslavsky,, 2015), and in particular, compressed representations of past input that let them predict future input (Shalizi and Crutchfield,, 2001; Bialek et al.,, 2006; Still,, 2014).
@@
Along these lines, Wolfram, (2023) makes a useful connection between LMs and the reducibility of computational processes. Some computational processes are chaotic and complex: there is no good way to systematically predict the digits of pi other than actually running the computation and learning what the next digit is. But other processes that appear very complex actually have learnable patterns that can be used to give simpler, compressed representations. For instance, $\frac{3227}{555}$ is $5.8144144144\dots$ where the $144$ repeats forever. This repeating structure means that we could give a compressed representation of what the millionth digit is without having to do additional computation. Wolfram, suggests that the success of LMs (which we know to be effective compressors) is proof that human language is amenable to this kind of compressed representation. A plausible candidate that allows for this compression is linguistic structure.
4.2 What LM experiments can in principle tell us about human language
@@
Armed with the framework of real patterns, we can think more critically about what LM experiments can tell us about linguistic theory and why we think that it is clearly possible to learn something about human language from LM experiments. Consider a thought experiment, this time using the linguistic concept of c-command (a particular relationship between elements in a linguistic tree which has been posited to be important for a variety of structure-sensitive linguistic operations). In this thought experiment, we get embedded representations for each word and do some fancy matrix multiplication between every pair of representations for each word in each sentence, to get a representation of the pairwise relationship between each pair of words in each sentence. We do some analysis on these word pair representations, and we find that, in layer $n$, neuron $m$ patterns in a very interesting way. If and only if the value of that particular neuron is positive, then the first word in the pair c-commands the second word. If it’s negative, then we know that it doesn’t. You experiment some more and find this is remarkably and perfectly consistent. (N.B.: We want to be very clear that this is a thought experiment; no experiment would ever give such clean results and LMs just don’t seem to represent anything in such linear and human-interpretable ways.)
@@
If we did discover a result like this, we would be justified in drawing several conclusions. First, the ease with which we extracted the c-command relationship would be incredibly compelling evidence that, for performing linguistic tasks in English, understanding which words c-command which other words is very useful exactly because, even though the model didn’t have anything about c-command built into it, it learned it emergently. Second, we would conclude that learning the c-command relationship is possible with relatively little built-in language-specific bias since a Transformer model learned it simply by optimizing for a word prediction objective. We would thus update our credence that the learning of complex syntactic relationships like c-command is possible without a built-in Universal Grammar (or that somehow the Transformer architecture is itself implementing a form of UG). Third, we could use this method to test various competing theories about which kinds of structural relationships should be more or less prominently represented in models. Perhaps most importantly, we could use these results to try to work out how an abstract structural relationship like c-command is represented in the weights of a system which, like the brain, is composed of cascades of activating neurons. Given our relatively impoverished understanding of how human neurons in the brain represent complex linguistic structures, such explorations could be informative as in principle in silico models of how the brain might process language.
It would be great if things were so easy! But they are not: the problem is that, like all complex systems, language models don’t learn and represent information in neat and interpretable ways where we can just read off what they are doing. (There is no reason for them to be interpretable; they are not optimized for it.) Rather, it takes a lot of work to ask questions like: are LMs representing part of speech (Tenney et al.,, 2019)? are they representing grammatical dependencies (Hewitt and Manning,, 2019; Manning et al.,, 2020; Chi et al.,, 2020)? how do they represent grammatical roles like subject? And the answers that come out are messy. A lot of this work has been limited by the best available techniques in interpretability. And, even if we did have perfect methods, the representations themselves would be messy and gradient, not neatly separable.
@@
What if we looked and looked and didn’t find the c-command relationship in our model representations? That’s a trickier scenario, since we might simply not have used the right methods to find it, and it’s still robustly represented. Or it might turn out that, while c-command is critical to human processing, LMs process language in some other way that has no recourse to c-command. Or it might turn out that neither humans nor models need to represent something like c-command to demonstrate proficient linguistic behavior. Even in this negative case, though, we think the results could be fruitful for generating hypotheses and stimulating further inquiry, grounded in data and requiring precise formulations of what exactly is entailed by testing for c-command.
@@
Thus, while the linguistics conclusions we can draw from LM intepretabilty experiments require care, we think it should be relatively uncontroversial that the idealized c-command experiment above would be linguistically interesting—even while recognizing a wide variety of opinions as to whether the current slate of results from LM experiments are linguistically interesting (Wilcox et al., 2023a, ; Mahowald et al.,, 2024; Lan et al.,, 2024).
@@
Our own stance is that, with improvements in interpretability techniques and particularly techniques focused on generating linguistic insights into the mechanistic structure of important phenomena l, these kinds of experiments are no longer mere pipe dreams. As Chris Potts said in a talk focused on LM processing of long-distance structures (drawing on Arora et al.,, 2024): “it seems to me that I can talk about the feature in either way, symbolic or neural, without any loss of accuracy” (Potts,, 2025). This broad line of work, exploring symbolic representations in systems that solve genuinely interesting linguistic tasks, has started to make good on the promise of Smolensky, (1988)’s prescient vision “in which traditional and connectionist theoretical constructs collaborate intimately to provide an understanding of cognition.”

### v2

@@ 3.3.2 The statistical tradition of linguistics
Perhaps the most well-known such contribution is distributional semantics: the idea that the semantics of a word is related to the distribution over contexts in which that word appears, often cited to Firth, (1957, p. 11), and developed more systematically by Harris, (1954). This idea originates from the school of structuralist linguistics (Saussure,, 1916; Bloomfield,, 1926). Structuralist linguistics had as one theoretical aim the development of discovery procedures, which were formal procedures that could be applied to bodies of text in order to discover (and even define) linguistic structures (Harris,, 1951). The most well-developed of these discovery procedures were statistical in nature. For example, Harris, (1955) developed a theory of words and morphemes based on statistical co-occurrence patterns, including a procedure for discovering morpheme boundaries by effectively calculating transitional probabilities, an idea taken up again much later in the psycholinguistics of language learning (Saffran et al.,, 1996), and closely related to tokenization methods such as byte-pair encoding (Shibata et al.,, 1999; Tanaka-Ishii,, 2021, Ch. 11). At issue was the relationship between grammatical structure and the observable statistical structure of a corpus of language—one of the core questions that linguists working on language models are interested in again today.
@@
Nevertheless, the statistical analysis of language continued under the heading of usage-based approaches (Bybee and Hopper,, 2001), in which the key to language is not a set of underlying formal rules but emergent properties based on the statistics and dynamics of language use (from which rules or rule-like behavior might emerge). Traditions in typological syntax (Greenberg,, 1963; Dryer,, 1992), functionalist syntax (Comrie,, 1989; Keenan and Comrie,, 1977), construction grammar (Croft,, 2001; Goldberg,, 2009), probabilistic modeling (Bresnan et al.,, 2001, 2007; Christiansen and Chater,, 2016), and evolutionary linguistics (Kirby and Hurford,, 2002) have all carried the banner of a thoroughgoingly statistical approach to language. Earlier arguments about the relationship between statistics and structure, and the value of probabilistic methods, mirror many of the points we are making here (Abney,, 1996; Pereira,, 2000; Manning,, 2003; Bresnan et al.,, 2007; Lappin and Shieber,, 2007; Norvig,, 2012).
@@
These statistical traditions of linguistics played a catalyzing role in the birth of neural networks and LMs. The connectionist framework from which neural LMs emerged overlapped with the statistical tradition in linguistics, often in conflict with the anti-statistical tradition. A major locus of this early work using neural networks was about the English past tense, with Rumelhart and McClelland, (1987) developing a neural model for forming the past tense from present tense words, triggering much debate (Pinker and Prince,, 1988). Elman, 1990a, a Linguistics PhD, developed an early recurrent neural network (RNN) sequence architecture () in order to solve the problem of finding linguistic structure in time, a problem motivated by findings in linguistics and psycholinguistics (Frazier and Fodor,, 1978). This basic architecture still underlies many language models (Feng et al., 2024a, ). Chris Manning, a pioneer in neural and probabilistic models, was trained as a linguist (Manning,, 1995), and argued for a statistical approach to syntax (Manning,, 2003), before going on to work on influential neural methods for performing linguistic tasks (e.g., Pennington et al.,, 2014).
@@
The history of the field and the dominance of the generative tradition made it such that these figures are sometimes seen as ‘not linguists’. But the traditions they emerged from were fundamentally concerned with questions about human language and cognition and belong under the heading of linguistics. To be clear, we do not think that the success of modern language models unequivocally supports or refutes any particular intellectual tradition in linguistics. But we do think that the narrow and exclusive theoretical focus of generative linguistics, coupled with its relative dominance within American linguistics departments throughout the late 20th century, caused tragic missed connections in intellectual history, and has left the field of linguistics diminished compared to where it could be.
@@
Progress in language science will not come from hyperfocus on any one goal, nor from any one theory of what language is, nor from one framework for understanding linguistic phenomena. Rather, the science of language needs to draw on a plurality of perspectives, from a wide variety of disciplines. In the current moment, that means leveraging the explosion of intellectual creativity springing forth around language models. Just as the concept of entropy arose from ad-hoc analysis of complex machines and ended up revolutionizing fundamental physics, it is possible and even likely that ideas based on language models will revolutionize linguistics.
@@ 4 Where does that leave the science of language?
If neural models make us rethink some aspects of linguistic theory, where do we go from here? What is the status of the formal structures discovered through linguistic analysis, if they don’t have to be innately latent in the human brain? We believe that linguistic structure is as real as it ever was, and that the LM revolution presents us with important new ideas and methods for studying language. Taking a statistical and information-theoretic approach to language brings these possibilities into focus. We also consider ways in which LMs might give us reason to update our linguistic theories to be more gradient, usage-based, and functionalist, by serving as a proof-of-concept system that implements ideas that were previously hard to formalize.
@@ 4.1 Linguistic structure is real
LMs are proof-of-concept that systems can process language without having linguistic structure hardwired in. But that doesn’t mean it isn’t real. We hold that linguistic structure is a real pattern in the sense of Dennett, (1989): it provides a compressed and useful representation of important aspects of language. To explain or describe linguistic phenomena without reference to linguistic structure, perhaps in terms of some more reductive neural theory, would be hopelessly complex — and would make understanding LM behavior more difficult (see Nefdt,, 2023, for a worked-out theory of linguistic structure as real patterns).
@@
Figure 4 shows a schematic of how this account works for a phenomenon like subject–verb agreement. Modern language models are proficient at this task, proficient enough that their behavior looks rule-like. But, given the dynamic and chaotic nature of neural networks, it is likely that there are messy and heterogeneous structures and processes that give rise to this rule-like behavior in models. These structures might only very noisily map onto linguistic categories like “grammatical subject” or “number agreement”. One reaction to that might be to conclude that linguistic theory was wrong or misguided—that it seemed like there was such a thing as “grammatical subject” that was crucial for verb agreement, but that actually the underlying reality was more complicated and now we can do away with all these rules and replace them with the messy, complicated neural system instead.
@@
This hyper-reductive conclusion would not be productive. A theory of language which points to complex neural circuits for explaining the data of subject–verb agreement, without referring to something about the concept of grammatical subject, would be like an example from Dennett, (1989, Ch. 2): imagine aliens who can perfectly predict human behavior using atomic physics (that clump of human atoms will move towards that coffee shop) but without understanding human beliefs and desires (that the human wants coffee and believes that the coffee shop will provide it). Lacking concepts of humans’ attitudes, the aliens’ theory of human behavior is perhaps perfectly accurate and reductive, but comes “at enormous cost of cumbersomeness, lack of generality, and unwanted detail” (Dennett,, 1990, p. 189). Dennett, concludes that the concept of belief is a real pattern in that it provides an abstraction that supports prediction and counterfactual reasoning based on coarse-grained data. The idea is that any abstraction that enables simple prediction, compression, and causal modeling in this way is a real pattern, fully deserving the epithet real even if there is a more reductive theory that lacks the abstraction. Similarly, even if it turned out we could explain linguistic behavior without recourse to anything recognizable as linguistic structure, it would come at the same cost of “cumbersomeness, lack of generality, and unwanted detail”.
@@
Figure 4: Schematic for linguistic structure as a real pattern, a (potentially leaky) abstraction that supports prediction and counterfactual reasoning based on coarse-grained data. In humans (in vivo), the complex physical, biological, and social processes associated with language learning, use, and knowledge may be abstracted into psycholinguistic theories describing operations on mental representations. In turn, the behaviors of these mental representations may be interpretable in terms of coherent linguistic structure. In neural networks (in silico), the complex patterns of weights and activations may be abstracted into interpretable circuits that approximately compute linguistically meaningful features like whether a word is singular and a grammatical subject. In turn, those interpretable circuits may be interpretable in terms of a larger abstraction of a coherent linguistic structure. Thus linguistic structure is a real pattern at the highest level of abstraction, which allows us to understand language as it is implemented by a human or by a neural network.
We join Dennett, in his conviction that real patterns are, well, real. As such, we are not eliminativists about grammatical structure, nor do we believe that this kind of eliminativism would be scientifically productive. Some in the generative tradition adopt a similar view, contra the idea that every posited linguistic structure must be in the head. Adger, (2022) writes: “Generative grammarians think of the posits of their theories in much the same way as physicists think of quarks or particles in wave form: They are the best explanations of the phenomena, though we do not know exactly what they correspond to in the world we can currently observe.” LMs are a promising avenue for figuring out what these explanations do, in fact, correspond to in the world.
@@
Indeed, the idea of real patterns can help explain the success of language models. Language models are trained to predict tokens in context. From an information-theoretic perspective, this means they are finding simple representations for linguistic input (Chater and Vitányi,, 2007), and in particular, compressed representations of past input that let them predict future input (Tishby et al.,, 1999; Shalizi and Crutchfield,, 2001; Bialek et al.,, 2006; Still,, 2014; Tishby and Zaslavsky,, 2015). Insofar as linguistic structure is real, it is part of the representation that enables this compression.
@@
A similar point is made by Wolfram, (2023), who connects the success of LMs with the reducibility of computational processes. Some computational processes are chaotic: there is no good way to systematically predict the digits of pi other than actually running the computation and learning what the next digit is. But other processes that appear complex have learnable patterns that can be used to give simpler, compressed representations. For instance, $\frac{3227}{555}$ is $5.8144144144\dots$ where the $144$ repeats forever. This repeating structure means that we could give a compressed representation of what the millionth digit is without having to do additional computation. Wolfram, suggests that the success of LMs proves that human language has structure that enables this kind of compression. This compressibility is responsible not only for LMs’ success, but for the fact that it is possible to construct linguistic theories. Compressibility is structure.
4.2 What LM interpretability can tell us about human language
We do not mean to suggest that linguists and cognitive scientists should identify compressible patterns in behavior, call it a real pattern, and declare victory. Perhaps the grand challenge of all of cognitive science is to uncover the internal mechanisms in neural networks and brains that underlie and implement behavioral patterns (Smolensky,, 1988). Finding these internal mechanisms is not a prerequisite for the reality of linguistic structure—the mechanisms may be irreducibly complex. But when mechanisms are found in neural networks, they can be informative not only about linguistic structure itself, but also about how language processing and learning may be implemented in the human brain.
Interpretability in an idealized thought experiment
@@ 4.2 What LM interpretability can tell us about human language
Consider a thought experiment, using the linguistic concept of c-command (a particular relationship between elements in a linguistic tree which has been posited to be important for a variety of structure-sensitive linguistic patterns). In this thought experiment, we get embedded representations for each word and use an interpretability method to get a representation of the pairwise relationship between each pair of words in each sentence. We analyze these word pair representations, and we find that, in layer $n$, neuron $m$ patterns in an interesting way. If and only if the value of that particular neuron is positive, then the first word in the pair c-commands the second word. If it’s negative, then we know that it doesn’t. Further experimentation reveals that this pattern is remarkably consistent. (N.B.: We want to be clear that no experiment would ever give such clean results; LMs just don’t represent anything in such human-interpretable ways.)
@@
If we did discover a result like this, we would be justified in drawing several conclusions. First, the ease with which we extracted the c-command relationship would be compelling evidence that, for performing linguistic tasks in English, understanding which words c-command which other words is useful — so useful that the model learned it emergently. Second, we would conclude that learning the c-command relationship is possible with relatively little built-in language-specific bias. We would thus update our credence that the learning of complex syntactic relationships like c-command is possible without a built-in Universal Grammar. Third, we could use this method to test various competing theories about which kinds of structural relationships should be more or less prominently represented in models. Fourth and most importantly, we could use these results to try to work out how an abstract structural relationship like c-command can be represented in the weights of any system which, like the brain, is composed of cascades of activating neurons. Such explorations would be informative as proof-of-concept in silico models of how the brain might process language.
@@
If we didn’t find the c-command relationship in our model representations, that’s a trickier scenario. We might simply not have used the right methods to find c-command, even though it’s robustly represented. Or it might turn out that, while c-command is critical to human processing, LMs process language sans c-command. Or it might turn out that neither humans nor models need to represent something like c-command to demonstrate proficient linguistic behavior. Even in this negative case, though, the results could be fruitful for generating hypotheses and stimulating further inquiry, grounded in data and requiring precise formulations.
@@
Thus, while the linguistic conclusions we can draw from LM interpretability experiments require care, it should be relatively uncontroversial that the idealized c-command experiment above would be linguistically interesting—even while recognizing a wide variety of opinions as to whether the current slate of results from LM experiments are linguistically interesting.
Interpretability in the real world
Our c-command example is not realistic: like all complex systems, language models don’t learn and represent information in neat and interpretable ways. Rather, it takes a lot of work to ask questions about how LMs represent parts of speech (Tenney et al.,, 2019) or grammatical dependencies (Hewitt and Manning,, 2019). And the answers that come out are messy.
But we are optimistic. The extent to which LMs are “blackboxes” is now overstated because of major advances in the field of neural network interpretability. For example, as discussed in Section 2.2, researchers now have a grip on how neural systems use the geometry of word embeddings to represent syntactic relations among words (Hewitt and Manning,, 2019; Chi et al.,, 2020; Eisape et al.,, 2022; Diego-Simón et al.,, 2024).
More recent techniques have shown, in a mechanistic way, exactly how artificial neural circuits can be used to perform higher-level computations (Lindsey et al.,, 2025), finding neural circuits and induction heads in Transformers (Elhage et al.,, 2021) or using sparse autoencoders to find key features in model representations (Huben et al.,, 2023). These techniques often depend on causal manipulation—showing that, if particular parts of the neural network are perturbed or patched in particular ways, the output is affected in predictable ways (Geiger et al.,, 2023). For instance, Wang et al., (2023) shows how it is possible to trace a circuit that controls the completion of an object (e.g., “When Mary and John went to the store, John gave a drink to ___.” where the intended answer is “Mary”). Many papers have now used these techniques to causally study what parts of a network are responsible for particular kinds of complex linguistic behavior like grammatical number agreement (Lasri et al.,, 2022; Finlayson et al.,, 2021; Mueller et al.,, 2022; Lakretz et al.,, 2019), verb conjugation (Hao and Linzen,, 2023), animacy processing (Hanna et al.,, 2023), and various kinds of long-distance dependencies (Arora et al.,, 2024). These techniques demonstrate how complex linguistic behavior can actually be implemented mechanistically in a neural system. These mechanisms are strong hypotheses for how syntactic relationships are represented in the human brain as well.
@@
It is our hope—and also our prediction—that interpretability techniques will make it possible to close the gap between linguistic theory and implementation (mechanistic interpretability is a big part of “the way forward” per Millière and Buckner,, 2024). This broad line of work, exploring symbolic representations in systems that solve genuinely interesting linguistic tasks, has started to make good on the promise of Smolensky,’s (1988) prescient vision “in which traditional and connectionist theoretical constructs collaborate intimately to provide an understanding of cognition”.